# Spatial-RAG: Spatial Retrieval Augmented Generation for Real-World Spatial Reasoning Questions

### v3

@@ 4.3.1 Sparse Spatial Relevance Scoring
where $g_{r}$ and $g_{t}$ are reference and target spatial objects, respectively. $d(g_{r},g_{t})$ is a distance function measuring proximity in the spatial database. If $g_{t}$ overlaps with $g_{r}$, we assign a perfect relevance score of 1.
This ensures that objects within a region are maximally relevant, while those outside the region receive scores that decay with increasing distance.
@@ 4.3.2 Dense Spatial Relevance Scoring
Extracting Spatial Requirements
Given a user query $q$ and a set of text descriptions $d_{t}$ for spatial objects $G_{t}$, we extract the relevant spatial content via an attention-based masking function:
@@
where $v_{q,s}$ and $v_{t,s}$ are dense vector representations of spatial features, and $\mathcal{M}$ is the extraction function mapping input text to a spatial related text, $\mathcal{E}$ is the text encoder.
Ranking via Cosine Similarity
The relevance score is computed via cosine similarity:
$f_{s}^{\text{dense}}=\frac{v_{q,s}\cdot v_{t,s}}{\parallel v_{q,s}\parallel\parallel v_{t,s}\parallel}.$ (11)
4.3.3 Hybrid Ranking as a Generalized Model
We can demonstrate that hybrid ranking generalizes both sparse and dense ranking approaches:
  • Sparse-Only Case: If $\lambda_{d}=0$ , then $f_{s}=\lambda_{s}f_{s}^{\text{sparse}}$, reducing to a purely distance-based ranking.
  • Dense-Only Case: If $\lambda_{s}=0$ , then $f_{s}=\lambda_{d}f_{s}^{\text{dense}}$, reducing to a purely semantic-based ranking.
  • Hybrid Case (General): If both weights are nonzero, hybrid ranking benefits from both explicit spatial constraints and implicit semantic relevance, leading to a more comprehensive ranking mechanism.
This formulation ensures that hybrid ranking outperforms any single-ranking approach by capturing both spatial proximity and semantic alignment.
@@ 4.4 Multi-objective Generation
The semantic candidate set $C_{k}$ and the semantic relevance score $f_{k}$ are calculated based on dense vector similarity, we put the details in appendix A. After all the scores and candidate sets are acquired, the problem becomes a multi-objective optimization problem since each perspective (spatial and semantic) contributes independently.
4.4.1 Pareto Front Computation
Given the spatial and semantic relevance scores, our goal is to identify the Pareto-optimal candidates that achieve the best trade-off between these objectives. A candidate $y$ is Pareto-optimal if no other candidate dominates it in both spatial and semantic relevance. Formally, the Pareto front $P(q)$ is defined as:
@@ 4.4.1 Pareto Front Computation
$f_{s}(q,y^{\prime})\geq f_{s}(q,y)\ \text{and}\ f_{k}(q,y^{\prime})\geq f_{k}(q,y),$
[…]
@@ 4.4.2 LLM-Based Trade-Off Decision
Once the Pareto front $P(q)$ is determined, we use an LLM to dynamically balance the trade-offs between spatial constraints and semantic preferences based on the context of the user query. Specifically, the LLM receives the user query, sparse spatial relevance scores, and spatial object descriptions as input:
@@
A dynamic weighting function $\lambda_{s},\lambda_{k}=h(I)$ based on contextual information is extracted from the input, adjusting the importance of spatial vs. semantic relevance, where $h$ is a learned function capturing query-specific trade-offs.
The top-ranked candidate $y^{*}$ is selected by LLM:
@@
and the LLM generates a natural language response.
The system adapts to different query contexts instead of using a fixed weighting scheme. By structuring decision-making into discrete steps (candidate filtering $\rightarrow$ Pareto selection $\rightarrow$ trade-off balancing $\rightarrow$ response generation), the LLM avoids generating infeasible or illogical results. This structured approach maximizes accuracy and usability, ensuring that the system’s final response aligns closely with the original user intent.
@@ 5.1 Experiment Setting
5.1.1 Datasets
@@
TourismQA (Contractor et al., 2019) dataset contains user questions crawled from TripAdvisor posts from 50 cities around the world. Reviews of restaurants, attractions, and hotels for each city are crawled from travel forums and hotel booking websites. We select two two popular tourist cities, New York City and Miami, to evaluate the performance of different methods. For New York City, the original dataset contains information on 9,470 Points of Interest (POIs), with a total of 17,448 QA pairs. For Miami, the dataset includes 2,640 POIs and 133 QA pairs. Our preprocessing steps on the dataset include removing POIs with empty review information and eliminating duplicate QA pairs.
5.1.2 Evaluation Metrics
We assess the experimental results from multiple dimensions: 1) Delivery Rate: Following (Xie et al., 2024), this metric assesses whether the method can deliver a result successfully. 2) Spatial Sparse Pass Rate: This metric evaluates whether the parsed spatial query is correct. 3) Spatial Dense Pass Rate: This metric evaluates whether the answer satisfies the spatial-related semantic constraints in the question. 4) Semantic Pass Rate: This metric evaluates whether the answer meets the semantic constraints in the question. We use GPT-4o to evaluate the last three metrics.
5.1.3 Models for comparison
City Model Delivery Rate Spatial Sparse Pass Rate Spatial Dense Pass Rate Semantic Pass Rate
NYC SD 100 - 57.1 32.1
TE 99.6 - 48.6 54.6
ST 100 - 53.6 55.9
Naive RAG 99.8 - 52.0 54.5
GeoLLM 99.8 - 69.5 42.8
Spatial-RAG (GPT-3.5-Turbo) 87.2 67.0 64.4 47.4
Spatial-RAG (GPT-4-Turbo) 86.1 65.0 71.6 50.1
Miami SD 100 - 36.8 22.4
TE 100 - 28.9 39.5
ST 100 - 26.3 42.1
Naive RAG 100 - 31.6 40.8
GeoLLM 100 - 52.6 36.8
Spatial-RAG (GPT-3.5-Turbo) 86.8 75.8 57.6 43.9
Spatial-RAG (GPT-4-Turbo) 86.8 81.8 51.5 45.5
Table 1: Performance comparison of models in New York City (NYC) and Miami. The framework is deployed on GPT-4-Turbo and GPT-3.5-Turbo and compared its performance with several baseline models, including SD, TE, ST, Naive RAG, and GeoLLM.
Model Delivery Rate Spatial Sparse Pass Rate Spatial Dense Pass Rate Semantic Pass Rate
Spatial-RAG (GPT-4-Turbo) 86.1 65.0 71.6 50.1
w/o sparse spatial 98.9 - 53.3 78.4
w/o dense spatial 89.3 61.5 68.4 49.6
w/o dense semantic 85.9 72.8 75.9 34.8
Table 2: Different modules of the proposed framework are removed for ablation study.
@@ 5.1.3 Models for comparison
We used GPT-3.5-Turbo and GPT-4-Turbo as the LLM in our framework to assess the impact of LLM capability on performance and compared it against the following four methods: Sort-by-distance (SD) (Contractor et al., 2021b): This method ranks the candidate spatial objects based on their distance to the reference objects in the spatial question. Text embedding (TE) (Cakaloglu et al., 2020): This greedy method minimizes the distance between the vector embeddings of the text description of the reference object and the target object. Spatial-text (ST): This approach computes the embeddings of the user’s question and compares the similarity between the question embedding and the text description embedding of the target object. Additionally, the object’s location is encoded as a distance score. The answer is then determined based on the weighted average of these scores. Naive RAG (Lewis et al., 2020): This method saves all spatial objects’ descriptions in a vector database and retrieves the most relevant objects based on vector similarity. GeoLLM (Manvi et al., 2024): This method encodes the spatial objects to address and enrich its context by adding spatial information of nearby spatial objects.
5.2 Main Results
The main results are summarized in 1 For the NYC dataset, Spatial-RAG demonstrates minimal differences in Delivery Rate when deployed with different LLMs. Approximately 86.1% of the questions are successfully processed. The failure cases where the framework is unable to generate valid recommendations can be classified into two main categories: Around 12.9% of the questions fail to retrieve any spatial objects from the spatial database, either due to polygon recognition errors or because the specified region in the SQL query lacks relevant spatial objects. The remaining 0.9% of cases stem from the LLM’s inability to correctly re-rank the retrieved results during the reranking process.
Additionally, the performance gap of our framework based on two LLMs in generating spatial SQL queries is insignificant. However, GPT-4-Turbo outperforms GPT-3.5-Turbo by approximately 7% in Spatial Dense Pass Rate and also demonstrates a slight advantage in Semantic Pass Rate, and both models perform slightly below the optimal values (by 5% and 8%, respectively). This is due to the trade-off made when jointly considering both spatial proximity and the user’s personal preferences.
Since the SD method returns the nearest spatial object, it achieves a higher spatial dense retrieval score than the TE, ST, and Naive RAG methods. However, its exclusive reliance on distance results in inferior performance on other metrics. For GeoLLM, as it considers only the names of spatial objects and their distances, it also achieves competitive performance in Spatial Dense Pass Rate. In contrast, both the TE and ST methods consider semantic contexts, leading to superior performance in this aspect compared to other models. The Naive RAG and ST methods jointly optimize for both spatial dense retrieval and user semantic dense retrieval, resulting in comparable performance between the two models.
For Miami, which differs significantly from NYC in terms of area and geography, Spatial-RAG also demonstrates strong performance. The performance of the baseline models almost remains consistent with the patterns observed in the previous analysis.
Figure 3: An example of how Spatial-RAG operates: Given a question, 1) Sparse Spatial Retrieval: an LLM parses the natural language question into a spatial SQL query for spatial database, retrieving spatial objects that meet the spatial constraints and the sparse spatial relevance scores. 2) Question Decomposition and Dense Retrieval: In parallel, Spatial-RAG decomposes the question into spatial and semantic components and compares them with the descriptions of spatial objects to perform dense retrieval, filtering out irrelevant ones. 3) LLM Reranking: A language agent balances the spatial and semantic aspects to rerank the candidate and generate the final answer.
@@ 5.3 Ablation Study
We conducted three ablation studies to assess the importance of different modules in our framework. We remove the sparse spatial module, dense spatial module, and dense semantic module, respectively, and denote the ablation versions as w/o sparse spatial, w/o dense spatial, and w/o dense semantic. The results are summarized in Table 2. It’s found that there is a significant increase in Delivery Rate after removing the sparse spatial module, as this eliminates the need to determine the geometry type. In this case, the model relies solely on semantic information, resulting in a lower spatial score but an increased emphasis on semantic relevance. Similarly, when the dense semantic module is removed, the Spatial Dense Pass Rate reaches its highest value, as the model depends entirely on spatial modules for retrieval. However, this also leads to a lower Semantic Pass Rate, as semantic information is no longer incorporated into the ranking process. Overall, our spatial-RAG framework integrates both spatial and semantic information, leveraging multi-objective optimization to provide more effective and well-balanced spatial object recommendations.
5.4 Case Study
An example of polyline search is given in Figure 3. Spatial-RAG effectively identifies the user’s intent within a noisy query, detecting that the user needs to travel between two locations (Yankee Stadium and Hilton New York) and is specifically looking for dinner recommendations (”any recos for dinner that evening - something simple”). Based on this, Spatial-RAG classifies the spatial operation as a ”point within polyline” query and applies a 500m buffer distance. Within this region, it filters restaurant spatial objects (shown in Fig. 4). In contrast, traditional methods, such as distance-based approaches, typically generate a buffer zone around a single point, which may not necessarily include POIs along the user’s travel route. Our framework provides a more context-aware understanding of the user’s spatial intent, enabling more precise and relevant recommendations based on both location and user preferences.
Figure 4: An example of a ”point within polyline” query. The user intends to travel between Yankee Stadium and Hilton New York, seeking dinner recommendations along the route. The buffered region (500m) is highlighted, and restaurants are selected.

### v4

@@ 4.3.1 Sparse spatial relevance scoring
where $g_{r}$ and $g_{t}$ are reference and target spatial objects, respectively. $d(g_{r},g_{t})$ is a distance function measuring proximity in the spatial database. If $g_{t}$ overlaps with $g_{r}$, we assign a perfect relevance score of 1. This ensures that objects within a region are maximally relevant, while those outside the region receive scores that decay with increasing distance.
@@ 4.3.2 Dense spatial relevance scoring
Extracting spatial requirements and ranking via cosine similarity
Given a user query $q$ and a set of text descriptions $d_{t}$ for spatial objects $G_{t}$, we extract the spatial requirements via an attention-based masking function:
@@
where $v_{q,s}$ and $v_{t,s}$ are dense vector representations of spatial features from query $q$ and text descriptions $d_{t}$. $\mathcal{M}_{s}$ is a function mapping input text to a spatial related text. $\mathcal{E}$ is the text encoder. The dense spatial relevance score $f_{s}^{\text{dense}}$ is computed via cosine similarity between two vectors.
4.3.3 Hybrid spatial scoring as a generalized model
We can demonstrate that hybrid spatial scoring generalizes both sparse and dense approaches:
  • Sparse-only case: If $\lambda_{d}=0$ , then $f_{s}=\lambda_{p}f_{s}^{\text{sparse}}$, reducing to a purely distance-based ranking.
  • Dense-only case: If $\lambda_{p}=0$ , then $f_{s}=\lambda_{d}f_{s}^{\text{dense}}$, reducing to a purely semantic-based ranking.
  • Hybrid case (General): If both weights are nonzero, hybrid spatial scoring benefits from both explicit constraints and implicit relevance.
This formulation ensures that hybrid spatial scoring outperforms any single-scoring approach.
@@ 4.4 Multi-objective generation
The semantic candidate set $C_{k}$ and the semantic relevance score $f_{k}$ are also calculated based on vector similarity, we put the details in Appendix A. After all the scores and candidate sets are acquired, the problem becomes a multi-objective optimization problem since each perspective (spatial and semantic) contributes from different aspects.
4.4.1 Spatial-semantic Pareto frontier computation
Given the spatial and semantic relevance scores, our goal is to identify the Pareto-optimal candidates that achieve the best trade-off between these objectives. A candidate $y$ is spatial-semantic Pareto-optimal if no other candidate dominates it in both spatial and semantic relevance. Formally, the Pareto frontier $P(q)$ is defined as:
@@ 4.4.1 Spatial-semantic Pareto frontier computation
$f_{s}(q,y^{\prime})\geq f_{s}(q,y)\ \text{and}\ f_{k}(q,y^{\prime})\geq f_{k}(q,y),\text{with at least one strict inequality}\}.$
@@ 4.4.2 LLM-based trade-Off decision
Once the Pareto frontier $P(q)$ is determined, we use an LLM to dynamically balance the trade-offs between spatial constraints and semantic preferences based on the context of the user query. Specifically, the LLM receives the user query $q$, sparse spatial relevance scores $f_{s}^{\text{sparse}}$, and spatial object descriptions $d_{y}$ as input:
@@
A dynamic weighting function $\lambda_{s},\lambda_{k}=h(I)$ based on contextual information is extracted from the input, adjusting the importance of spatial vs. semantic relevance, where $h$ is a learned function capturing query-specific trade-offs. The top-ranked candidate $y^{*}$ is selected by LLM:
@@
and the LLM generates a natural language response. The system adapts to different query contexts instead of using a fixed weighting scheme. By structuring decision-making into discrete steps (candidate filtering $\rightarrow$ Pareto selection $\rightarrow$ trade-off balancing $\rightarrow$ response generation), the LLM avoids generating infeasible or illogical results.
@@ 5.1 Experiment setting
Table 1: Overview of Datasets
Dataset #POIs #QA Pairs
TourismQA-NYC 9,470 17,448
TourismQA-Miami 2,640 133
MapQA-ADJ 92,415 50
MapQA-AME 92,415 231
Datasets.
@@
We conduct our experiments on four datasets, including TourismQA-NYC Dataset and TourismQA-Miami Dataset datasets, which contains user questions about points of interest (POIs) in real-world cities. The questions are crawled from TripAdvisor posts, while the reviews of restaurants, attractions, and hotels are collected from travel forums and hotel booking websites. For both datasets, we apply similar preprocessing steps, including removing POIs with missing review information and eliminating duplicate QA pairs. MapQA-Adjacent Dataset (MapQA-ADJ) features questions such as “What [amenity] is adjacent to [location]?”, emphasizing topological relationships. MapQA-Amenities_Around_Specific Dataset (MapQA-AME) focuses on proximity-based queries like “What are the [amenity] within 50m of [location]?” The number of POIs and QA pairs for each dataset is summarized in Table 1.
Evaluation metrics.
To assess the recommendation performance of different methods, we evaluate each model using five widely adopted metrics: Precision@k (The proportion of the top-k returned items that are relevant to the query), Recall@k (The proportion of all relevant items for a query that appear within the top-k results), F1@k (The harmonic mean of Precision@k and Recall@k), and NDCG@k (Normalized Discounted Cumulative Gain up to rank k), where $k\in\{1,3,5,10\}$.
Table 2: Performance comparison of models in TourismQA-NYC and TourismQA-Miami.
Dataset Method Precision Recall F1 NDCG
@1 @3 @5 @10 @1 @3 @5 @10 @1 @3 @5 @10 @1 @3 @5 @10
TourismQA-Miami GeoLLM 0.3158 0.2719 0.2368 0.2053 0.0365 0.0945 0.1246 0.1824 0.0591 0.1094 0.1311 0.1623 0.3158 0.3200 0.3268 0.3600
Naive RAG 0.2973 0.2252 0.2054 0.1973 0.0814 0.1124 0.1491 0.2301 0.1005 0.1099 0.1310 0.1642 0.2973 0.2753 0.2849 0.3403
TE 0.3421 0.3070 0.2316 0.1842 0.0567 0.1544 0.1821 0.2231 0.0781 0.1558 0.1515 0.1543 0.3421 0.3766 0.3890 0.4748
SD 0.1053 0.0351 0.0211 0.0105 0.0075 0.0075 0.0075 0.0075 0.0138 0.0119 0.0105 0.0082 0.1053 0.1053 0.1053 0.1053
ST 0.3947 0.1316 0.0789 0.0395 0.0873 0.0873 0.0873 0.0873 0.1115 0.0730 0.0569 0.0386 0.3947 0.3947 0.3947 0.3947
Spatial-RAG(GPT3.5-Turbo) 0.5455 0.4141 0.4000 0.3152 0.1081 0.2072 0.2972 0.3732 0.1540 0.2262 0.2765 0.2817 0.5455 0.492 0.5276 0.5440
Spatial-RAG (GPT4-Turbo) 0.5152 0.4242 0.3758 0.2939 0.1026 0.2037 0.2841 0.3565 0.1454 0.2246 0.2619 0.2656 0.5152 0.4852 0.4983 0.5079
TourismQA-NYC GeoLLM 0.3650 0.3292 0.3020 0.2725 0.0168 0.0433 0.0614 0.1033 0.0311 0.0688 0.0907 0.1328 0.3650 0.3626 0.3708 0.4248
Naive RAG 0.4923 0.4447 0.4245 0.4181 0.0214 0.0556 0.0889 0.1678 0.0399 0.0932 0.1345 0.2167 0.4923 0.4551 0.4528 0.5058
TE 0.2250 0.2433 0.2395 0.2325 0.0101 0.0330 0.0506 0.0934 0.0189 0.0541 0.0767 0.1199 0.2250 0.2873 0.3533 0.5142
SD 0.2425 0.2492 0.2455 0.2402 0.0103 0.0314 0.0520 0.0945 0.0193 0.0520 0.0758 0.1194 0.2425 0.2694 0.2926 0.3624
ST 0.4725 0.4608 0.4460 0.4323 0.0238 0.0644 0.1026 0.1846 0.0433 0.1026 0.1475 0.2281 0.4725 0.4686 0.4745 0.5296
Spatial-RAG (GPT3.5-Turbo) 0.4611 0.4352 0.4144 0.4098 0.0188 0.0485 0.0751 0.1493 0.0350 0.0828 0.1188 0.2000 0.4611 0.4545 0.4534 0.5097
Spatial-RAG (GPT4-Turbo) 0.5665 0.4875 0.4555 0.4251 0.0274 0.0611 0.0908 0.1691 0.0483 0.0996 0.1384 0.2153 0.5665 0.5174 0.5065 0.5574
Table 3: Performance comparison of models on MapQA-ADJ and MapQA-AME.
Dataset Metric Methods
SD TE ST Naive RAG GeoLLM Spatial-RAG (GPT-3.5-Turbo) Spatial-RAG (GPT-4.0-Turbo)
MapQA-ADJ Precision 0.4600 0.4200 0.5600 0.3455 0.4558 0.5467 0.5057
Recall 0.4200 0.3750 0.5150 0.6170 0.6685 0.5350 0.7819
F1 0.4333 0.3880 0.5280 0.4067 0.5073 0.5247 0.5625
NDCG 0.4600 0.4200 0.5600 0.5866 0.6132 0.5800 0.7391
MapQA-AME Precision 0.4323 0.2751 0.5714 0.2452 0.3729 0.5566 0.6645
Recall 0.3923 0.2351 0.5173 0.6075 0.8333 0.7446 0.9072
F1 0.4054 0.2482 0.5346 0.3263 0.4744 0.6100 0.7298
NDCG 0.4323 0.2751 0.5714 0.6018 0.8199 0.7264 0.8719
Table 4: Ablation results.
Method Precision Recall F1 NDCG
@1 @3 @5 @10 @1 @3 @5 @10 @1 @3 @5 @10 @1 @3 @5 @10
Spatial-RAG(GPT4-Turbo) 0.5665 0.4875 0.4555 0.4251 0.0274 0.0611 0.0908 0.1691 0.0483 0.0996 0.1384 0.2153 0.5665 0.5174 0.5065 0.5574
w/o sparse spatial 0.3807 0.3545 0.3503 0.3307 0.0147 0.0431 0.0701 0.1305 0.0276 0.0701 0.1058 0.1667 0.3807 0.3608 0.3570 0.3455
w/o dense spatial 0.5569 0.4954 0.4628 0.4295 0.0263 0.0606 0.0919 0.1679 0.0466 0.0986 0.1392 0.2132 0.5569 0.5240 0.5132 0.5587
w/o dense semantic 0.4986 0.4311 0.4165 0.3806 0.0240 0.0553 0.0844 0.1499 0.0421 0.0883 0.1262 0.1896 0.4986 0.4679 0.4739 0.5219
Scratch 0.5392 0.4935 0.4706 0.4343 0.0236 0.0535 0.0829 0.1574 0.0435 0.0916 0.1323 0.2105 0.5392 0.5105 0.5042 0.5374
w/o RAG 0.3584 0.3333 0.3028 0.2644 0.0121 0.0323 0.0484 0.0846 0.0231 0.0568 0.0796 0.1191 0.3584 0.3570 0.3653 0.4186
Models for comparison.
@@
We used GPT-3.5-Turbo and GPT-4-Turbo as the LLM in our framework and compared them against the following methods: GeoLLM manvigeollm encodes the spatial objects to address and enrich their context by adding spatial information of nearby spatial objects. Naive RAG lewis2020retrieval saves all spatial objects’ descriptions in a vector database and retrieves the most relevant objects based on vector similarity. Text embedding (TE) cakaloglu2020text is a greedy method minimizing the distance between the vector embeddings of the text description of the reference object and the target object. Sort-by-distance (SD) contractor2021joint ranks the candidate spatial objects based on their distance to the reference objects in the spatial question. Spatial-text (ST) computes the embeddings of the user’s question and compares the similarity between the question embedding and the text description embedding of the target object. Additionally, the object’s location is encoded as a distance score. The answer is determined based on the average of these scores.
5.2 Spatial-RAG vs. Baselines
The comparative performance results across methods and datasets are summarized in Tables 2 and 3. Our Spatial-RAG (GPT4-Turbo) consistently surpasses baselines, including Naive RAG, GeoLLM, and template-based methods (ST, TE, SD). Specifically, a 19.9% improvement in Precision@1 was achieved by our Spatial-RAG (GPT-4-Turbo) compared to the best baseline model (ST) on TourismQA-NYC. It improved NDCG by 32.0% on MapQA-ADJ and by 52.6% on MapQA-AME compared to ST. On the MapQA-AME dataset, it notably increased Recall by 75.4% over ST. Compared to GeoLLM, Spatial-RAG (GPT-4-Turbo) achieved 95.4% higher Recall@10 on TourismQA-Miami. This advantage arises primarily from Spatial-RAG’s structured spatial retrieval pipeline, which leverages a spatial database to accurately retrieve relevant candidates that satisfy the spatial constraints, a capability lacking in embedding-based methods like Naive RAG and purely textual or parametric models (GeoLLM, SD, TE). Meanwhile, although ST occasionally leads in single-relationship precision (e.g., Precision@1 on MapQA-ADJ), Spatial-RAG’s GPT4-powered re-ranking provides superior overall performance through nuanced multi-constraint spatial reasoning. Among evaluation metrics, Spatial-RAG sees greater gains in Recall and NDCG than in Precision, highlighting that structured spatial queries significantly enhance completeness of retrieved results and the accuracy of ranking order. Finally, Spatial-RAG (GPT4-Turbo) consistently outperforms its GPT3.5 counterpart, underscoring GPT4’s advanced reasoning capabilities and improved candidate-ranking accuracy.
Replacing GPT3.5-Turbo with GPT4-Turbo within the identical retrieval pipeline contributes a further +2 pp on average, indicating that model capacity and spatially constrained retrieval are additive rather than redundant. This improvement is especially evident on structured spatial QA tasks such as those in MapQA, which feature well-defined relations like “adjacent” and “within 50m.” In contrast, the performance gap is narrower on noisier, open-domain datasets like TourismQA, suggesting that GPT4’s capacity brings greater benefit when precise spatial reasoning is required. The above results establish Spatial-RAG as the new state of the art, delivering statistically robust, across-the-board improvements relative to all prior baselines.
@@ 5.3 Ablation study
To validate our technical contributions, we conducted five ablation studies. First, we examined the impact of removing individual modules: the sparse spatial module (w/o sparse spatial), dense spatial module (w/o dense spatial), and dense semantic module (w/o dense semantic). Next, we evaluated two system-level variations: generating SQL from scratch (Scratch) and removing the RAG component (w/o RAG).
Table 4 summarizes the ablation results. Removing any component leads to a performance decline, highlighting the importance of each module. Notably, excluding the sparse spatial component results in the most significant drop, underscoring the value of integrating the spatial database with the LLM. The Scratch setting shows only a slight decrease, suggesting that GPT4-Turbo can internally formulate SQL queries when given essential inputs, even without templates. The w/o RAG variant performs worst across all metrics, indicating that retrieval-augmented generation is fundamental to the system’s overall effectiveness.
5.4 Case studies
Figure 3: Three POI queries with different reference objects.
The visualizations of three POI queries with different reference objects are given in Figure 3. Spatial-RAG effectively identifies the user’s intent within noisy queries, detecting that the user needs to travel near a location, along a route, or within a region. Based on this, Spatial-RAG applies proper spatial constraints to filter spatial objects. In contrast, traditional methods, such as distance-based approaches, always generate a buffer zone around a single location, which may not necessarily include POIs along the user’s travel route. Our framework provides a more context-aware understanding of the user’s spatial intent, enabling more precise and relevant recommendations based on both location and user preferences.